HUBBLE SPACE TELESCOPE DAILY REPORT #1015 PERIOD COVERED: 0800 EST Daily Status Report as of 319/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of PC 5072 (WF/PC Cycle 3 Calibration: PC Flats Wide. 1) Both the Wide Field Camera and the Planetary camera were used to make Earth Flat calibration exposures using the wide and medium band filters. 12 observations were received and displayed in the OSS. Three exposures were underexposed, six were saturated, and three were reported to be okay. 1.2 Completed Two Sets of PC 4929 (WF/PC Cycle 2: Improved Dark Calibration) The Planetary Camera was used to make two dark calibration images. The two exposures lasted 1200 seconds each, and the camera was pre-flashed before the observations. Both observations were reported to be normal exposures. 1.3 Completed Two Sets of PC 5014 (WF/PC Cycle 3: Calibration: PC Flats Narrow. 2) The Planetary Camera was used to make two standard Earth calibrations images for flat fielding of PC observations. Medium and narrow filters were used, with and without neutral density filters. The observations were executed as scheduled. Both exposures were reported to be underexposed. 1.4 Completed GHRS/2 3755 (The Absorption Cross Sections of Nearby Galaxies) The High Resolution Spectrograph was used to study the absorption line spectra produced by the halo of nearby galaxy 0026+129 by using the continuum of a background quasar as a standard source. Of particular interest is the spectra lines of Mg II and Carbon IV. The onboard target acquisition was successful. No anomalies were reported. 1.5 Completed FOS/BL/RD 5049 (FOS Cycle 3 Wavelength Calibration: Internal/ External Offsets) The Faint Object Spectrograph was used to make observations of external object HD207757 and also of the internal lamp to determine the offset for pixel to wavelength calibration. Object HD207757 was used because it lacks sufficient emission lines for grating G130H. The observations were executed as scheduled, and no problems were reported. 1.6 Completed WFC 5075 (WF/PC Cycle 3 Calibration: WFC Flats Wide. 1) Observations were made to make a standard calibration to obtain Earth flats using the Wide Field Camera. The observations were taken. No anomalies were reported. 1.7 Completed Two Sets of WFC 5076 (WF/PC Cycle 3 Calibration: WFC Flats Wide. 2) Routine Earth flat calibration exposures were taken with the Wide Field Camera. The observations were monitored in the OSS, and the data quality was assessed. No problems were reported. 1.8 Completed GHRS/2 5230 (GHRS Side 2 Carrousel Functional Test - Cycle 3) The High Resolution Spectrograph was used to repeat the orbital verification carrousel functional test. The engineering telemetry will be compared to determine if the carrousel response has changed in any way. The OSS monitored the Observations and assessed the data quality. No anomalies were reported. 1.9 Completed WFC 5023 (Cal/WFPC Cycle 3 Polarization Calibration) The Wide Field Camera was used to make polarization calibration observations by using images of the non-polarized standard star GD319. This will allow for the absolute transmittance of the polarized optical path to be established. Four images were taken, two were received in real-time and were reported to be okay. The other two observations still remain on the science tape recorder, and will be analyzed later. 1.10 Completed FOS/BL 5078 (Spectroscopic Observations of the Exposed White Dwarfs in the Dwarf Nova) The High Resolution Spectrograph was used to make four observations of the dwarf nova VW Hyi, which is normally at 13.6 magnitude, but can get up to 8th magnitude in outbursts. Four observations were received and displayed. There was a recentering event which occurred during one of the observations. The other observations were reported to look normal. 1.11 Completed GHRS/2 3706 (The Nature of Gaseous Loops in the Milky Way Halo) The High Resolution spectrograph was used to make spectral observations of star HD121800 which is located at high galactic latitude. The line of sight of this star passes through neutral hydrogen clouds which are moving toward the plane of the galaxy. This proposal called for the Return-to-Brightest onboard target acquisition. Instead, the acquisition stopped as soon as it detected some light from the Point Spread Function. As a result, the target was almost certainly outside the aperture during the science observation. 1.12 Completed FOS/RD 4528 (SINS: The Supernova Intensive Study - SN1987) The Faint Object Camera (f/96) was used to make observations of the supernova SN1987 which is located in the Large Magellanic Cloud. This is a continuation of proposal 3853. This is to obtain images of the remnant with several filters over a long period of time to obtain photometric data. The guide star acquisition did not complete in time for the shutter to open. The first three observation was executed with the shutter closed. The OSS successfully uplinked a command to open the shutter, and the last two observations were reported to look fine. 1.13 Completed HSP/UV2 5256 (A Search for UV Variability in GQ Muscae - Rpt. for HOPR#141) The High speed Photometer was used to make ultraviolet observations of the variable star GQ MUS (aka, Nova MUS 1983) to study its physical conditions near the white dwarfs in the system. Bypass filter F140LP was used. The onboard target acquisition was successful and showed the target well centered. The science observations were reported to be okay. There were two recentering events which occurred during the observations. 1.14 Completed FOS/BL 3441 (Post Asymptotic Giant Branch Evolution in the Magellanic Clouds) The Faint Object Spectrograph was used to make three spectral observations of the planetary nebula LMC-SMP-47 which is located in the Large Magellanic Cloud. The observations were reported to be well centered and to look okay. 1.15 Completed GHRS/2 5040 (GHRS Detector Calibrations - Cycle 3) The High Resolution spectrograph was used to make a set of internal calibrations to check the stability of the HRS detectors. The test included dark counts, focus checks. and flat field observations. The observation was executed as scheduled. 11 observations were received and displayed in the OSS. They were reported to look satisfactory. 1.16 Completed PC 4308 (Snapshot Survey For Companions to Planetary Nebula Nuclei - Cycle 3 High) The Planetary Camera was used to make an observation of the central region of planetary nebula IC4997. It is possible that the central star in a double star. The observations was received and displayed in the OSS, the data was reported to look okay. 1.17 Completed FGS 4893 (Orbital Motions of Close Binary Stars - Cycle 3 High) Fine Guidance Sensor #3 was used to make astrometric observations of the binary star system HD73900 to obtain its orbital elements. These observations are important since most of the estimates of stellar mass are made by observing elements of these systems. The astrometry was monitored and no problems were seen in the telemetry. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 13 Successful : 13 Scheduled Reacquisitions : 35 Successful : 35 2.2 FHST Updates: Scheduled : 21 Successful : 21 2.3 Significant HSTARs: None 2.4 Operations Note: A partial eclipse of the sun occurred at 317/1959Z through 2009ZZ on Friday Nov. 12,1993, and special eclipse management commands were successfully uplinked to the spacecraft. No anomalies were reported. JIS #5 (Flight Day 3 Rendezvous) Simulation was initiated this morning, Monday, Nov. 15, 1993 starting at 9 AM EST. It is scheduled to last for 10 hours. 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete WFC 4929 (WF/PC Cycle 2: Improved Dark Calibration) 3.2 Complete FOS/RD 4817 (Quasar Absorption Line Survey: Cycle 2 Observations) 3.3 Complete Three Sets of WFC 5074 (WF/PC Cycle 3 Calibration: WFC Flats Narrow. 1) 3.4 Complete Three Sets of WFC 4381 (HST Medium - Deep Survey - Cycle 3) 3.5 Complete FOS/RD 3837 (Spectropolarimetry of Low Red-Shift Active Galactic Nuclei) 3.6 Complete WFC 4658 (Determination of the Extragalactic Distance Scale. III M-101) 3.7 Complete PC 4308 (Snapshot Survey for Companions to Planetary Nebula Nuclei - Cycle 3 High) 3.8 Complete FOS/BL 3982 (Physical Conditions in the Narrow-Lined Region) 3.9 Complete WFC 5013 (WF/PC Cycle 3 Calibration: WFC Flats Narrow. 2) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-